# Achieving more human brain-like vision via human EEG representational alignment Zitong Lu 1,*, Yile Wang 2, Julie D. Golomb1

### v2

@@ Abstract
Despite advancements in artificial intelligence, object recognition models still lag behind in emulating visual information processing in human brains. Recent studies have highlighted the potential of using neural data to mimic brain processing; however, these often rely on invasive neural recordings from non-human subjects, leaving a critical gap in understanding human visual perception. Addressing this gap, we present, for the first time, ‘Re(presentational)Al(ignment)net’, a vision model aligned with human brain activity based on non-invasive EEG, demonstrating a significantly higher similarity to human brain representations. Our innovative image-to-brain multi-layer encoding framework advances human neural alignment by optimizing multiple model layers and enabling the model to efficiently learn and mimic human brain’s visual representational patterns across object categories and different modalities. Our findings suggest that ReAlnet represents a breakthrough in bridging the gap between artificial and human vision, and paving the way for more brain-like artificial intelligence systems.
@@ 1 Introduction
While current vision models in artificial intelligence (AI) are advanced, they still fall short of capturing the full complexity and adaptability inherent in the human brain’s information processing. Deep convolutional neural networks (DCNNs) have reached a performance level in object recognition that rivals human capabilities [1], and many studies have identified representational similarities in the hierarchical structure between DCNNs and the ventral visual stream [2, 3, 4, 5, 6]. However, the current alignment between DCNNs and human neural representations, while promising, still presents significant opportunities for further exploration and enhancement. Enhancing the resemblance between visual models and the human brain has become a critical concern for both computer scientists and neuroscientists. From a computer vision perspective, brain-inspired models often exhibit higher robustness and generalization, crucial for realizing true brain-like intelligence; meanwhile, from a cognitive neuroscience perspective, models that more closely mirror brain representations can significantly aid in our exploration of the brain’s visual processing mechanisms.
@@
Given these challenges and limitations, the pivotal question arises is how we can leverage our understanding of the human brain to enhance current AI vision models. Conventional approaches have limitations in emulating the complexity of the human brain’s visual information processing, even with increased model depth and layers [7]. This limitation has prompted the exploration of new methodologies. Researchers have attempted various strategies, including altering the model’s architecture (adding recurrent structures [8, 4, 9, 10, 11], dual-pathway models [12, 13, 14, 15, 16], topographic constraints [17, 18, 19, 20] or feedback pathways [21] ) and changing the training task (using self-supervised training [22, 23] or 3D task models [24]). However, can we directly use human neural activity to align ANNs on object recognition and achieve more human brain-like vision models?
@@
Several previous studies have begun to try to apply neural data to machine learning especially deep learning models. The earliest attempt was to apply human fMRI signals to amend the classification boundary of SVMs and CNNs to achieve better category classification performance [25]. Some more recent studies started to let the models learn neural representations. One common way is to add a similarity loss to increase the representational similarity between models and neural activity (neural recordings obtained from mouse V1, monkey V1 or IT) during the training [26, 27, 28, 29]. Another strategy from [30] is to add an additional task based on an encoding module to predict monkey V1 neural activity. Both similarity-based method and multi-task framework can achieve more brain-like representations and improve model robustness. However, these neural alignment studies have two key challenges: (a) Dependence on animal instead of human neural activity. This limits the direct applicability and relevance of findings to human visual processing, and it is harder to enable models to effectively learn the human brain’s representational patterns because human non-invasive recording data usually have significantly lower data quality than data from animal invasive recordings. (b) Single brain region or single model layer alignment. On the one hand, previous studies just aligned a single early or late layer in CNN and/or align the model with a certain brain region, V1 or IT. On the other hand, it remains unclear which specific brain region should align with which particular layer of the model, leading to potential misalignment and inaccuracies. Additionally, a recent study focused on video emotion recognition first applied a representational similarity-based method to align CNN with human fMRI activity [31]. However, it is noteworthy that they focused on 6-category emotion classification tasks, may fall short in the more complex and diverse domain of object recognition which has larger space and multitude of object categories. Therefore, our work addressed this by employing an additional encoding module that goes beyond mere similarity. This module predicts human neural activity and is trained to autonomously extract complex visual features, offering a more effective approach for aligning the model with human neural representations in object recognition.
To bridge the gap between AI vision and human vision, we propose a more human brain-like vision model, ReAlnet, effectively aligned with human brain representations obtained from EEG recordings, based on a novel and effective encoding-based multi-layer alignment framework. Our representational alignment framework allows us to obtain personalized vision models by aligning with individual’s neural data. To the best of our knowledge, we are the first to directly align object recognition models using non-invasive neural data recorded from human brains, which opens new possibilities for enhancing brain-like representations in models based on human brain activity. Moreover, the human brain-aligned ReAlnet shows improved similarity to human brain representations across different modalities (both human EEG and fMRI) and human behaviors.
@@ 2.1 Aligning CORnet with human EEG representations
We applied a novel image-to-brain multiple-layer encoding alignment framework which let the model not only accurately classify the object category but also generate the realistic EEG signals via minimizing both classfication and generation losses during the training (Figure 1A, see Methods). Based on this alignment framework, we build ten individual ReAlnets, using the state-of-the-art CORnet-S model [32, 9] as the foundational architecture. Each ReAlnet, which has the same architecture as CORnet, is additionally trained on a real human subject’s EEG signals, recorded while viewing a massive number of natural images from THINGS EEG2 [33] training set. After training, We employed an independent test dataset consisting of 200 images and associated EEG activity from the THINGS EEG2 test set. These test set images had not been presented at all during the training process, coming from entirely novel (untrained) object categories. For models (ReAlnet and COrnet), we input these 200 images to each model and obtain the feature vectors corresponding to each image for each layer in the model. Then we calculated the temporal similarity between different models and human brain EEG based on the representational similarity analysis (RSA) method. Figure 1B shows the representational similarity between different models and temporal human brain EEG signals for early and late layers in ReAlnet, CORnet-S, ResNet101, and CLIP (with ResNet101 backbone) respectively, and ReAlnet aligned with human neural representations shows the highest similarity to the human brain.
@@
In this study, our core focus is to investigate whether aligning the model with individual neural representations of humans can enhance the model’s similarity to the human brain. Our assessment of the models’ similarity to humans is not limited to its similarity with human brain EEG representations (based on THINGS EEG2 test set [33]); Similar to EEG, we then evaluated the model’s similarity to human brain fMRI representations (a completely different modality) from human subjects viewing novel image categories (based on Shen fMRI test set [34]). Additionally, we measured the similarity between the model and human behavior in several object recognition tasks using the Brain-Score platform [35] based on two behavioral benchmarks. For methodological details, refer to our Methods section. In the following sections, we elaborate on how ReAlnet – compared with the original CORnet not aligned with human neural data – more closely resembles human brain representations.
@@ 2.2 Improved similarity to human EEG
Here, for each of the 10 human subjects from the THINGS EEG2 dataset [33], we calculated (1) the similarity between their EEG data and CORnet (with the same structure as ReAlnet, but non aligned human neural data and non-individualized model), and (2) the similarity between their EEG data and the subject-matched ReAlnet via representational similarity analysis (RSA) based on THINGS EEG2 test dataset (See more details in Methods section). ReAlnets show significantly higher similarity to human EEG neural dynamics for all four visual layers (Layer V1: 70-130ms and 160-200ms; Layer V2: 60-200ms; Layer V4: 60-200ms; Layer IT: 70-160ms) than the original CORnet without human neural alignment (Figure 2A). Further statistical analysis of each layer’s similarity improvement (ReAlnet - CORnet) and improvement ratio ((ReAlnet - CORnet) / CORnet) also indicate that at the similarity peak timepoint, there is a maximum of an 8% similarity improvement and an 80% improvement ratio, with the average improvement for the 50-200ms time-window being over 5% and the average improvement ratio over 40% (Figure 2B). It is worth noting that the test set doesn’t overlap with the training set in terms of object categories (concepts). Therefore, these significant improvements reveal ReAlnet’s generalization capability across different object categories.
@@
Additionally, unlike traditional models in computer vision, ReAlnet is a personalized model trained based on different individuals’ neural data. This sparked our interest in exploring whether these personalized ReAlnets exhibit intra-model individual variabilities and how such variabilities change across different layers of the model. To investigate this, we conducted comparisons between model RDMs based on 200 images in THINGS EEG2 test set across different layers, using the dissimilarity (one minus the Spearman correlation coefficient) between two RDMs corresponding to two ReAlnets as an individual variability index. And our results show: (1) Personalized ReAlnets indeed exhibit individual variability (Figure 2C). (2) This variability increases with the depth of the layers (from Layer V1 to Layer IT, Figure 2C). This may also suggest a trend of increasing individual variability from primary to higher visual cortical areas in human brains.
@@ 2.3 Improved similarity in ReAlnets to human fMRI
Although ReAlnet demonstrates higher similarity to human EEG, a question arises: Does ReAlnet learn representations specific to EEG, or more general neural representations of the human brain? To ensure that our alignment framework enables the model to learn representations beyond the single modality of EEG, we utilized additional human fMRI data of three human subjects viewing natural images to evaluate the model’s cross-modality representational similarity to human fMRI.
Excitingly, we indeed observed an increase in this cross-modal brain-like similarity. Based on human fMRI signals of three subjects viewing 50 natural images, the similarity results indicate that even though ReAlnets were aligned based on human EEG data from different participants, they still resemble the human brain more closely on fMRI data compared to CORnet (Figure 3A). Also, there is a significant correlation of ReAlnets’ similarity improvement compared to CORnet between EEG and fMRI () (Figure S1).
@@
Additionally, to further assess the enhanced similarity between ReAlnet and human brain representations during image observation - not solely with natural images - we also measure the similarity between models and human fMRI signals from three subjects viewing 40 artificial shape and 10 alphabetical letter images in Shen fMRI test dataset. Although these images might be outside the natural image distribution, our results further demonstrate ReAlnet’s improved similarity to human brain representations in comparisons to CORnet (Figure 3B-C).
@@
These findings further highlight three points: (1) Across multiple ROIs, ReAlnets exhibits higher human fMRI similarity than CORnet. (2) Despite not being trained with the EEG data of subjects in the fMRI dataset, almost every ReAlnet shows higher fMRI similarity, suggesting that ReAlnet learns consistent brain information processing patterns across subjects. (3) Images from fMRI dataset for evaluation were never presented during the alignment training, reaffirming the generalization of ReAlnets in improving brain-like similarity across object categories and images.
2.4 Improved similarity in ReAlnets to human behavior

### v3

# Achieving more human brain-like vision via human EEG representational alignment Zitong Lu 1,2,*, Yile Wang 3, Julie D. Golomb1

@@ Abstract
Despite advancements in artificial intelligence, object recognition models still lag behind in emulating visual information processing in human brains. Recent studies have highlighted the potential of using neural data to mimic brain processing; however, these often rely on invasive neural recordings from non-human subjects, leaving a critical gap in understanding human visual perception. Addressing this gap, we present, ‘Re(presentational)Al(ignment)net’, a vision model aligned with human brain activity based on non-invasive EEG, demonstrating a significantly higher similarity to human brain representations. Our innovative image-to-brain multi-layer encoding framework advances human neural alignment by optimizing multiple model layers and enabling the model to efficiently learn and mimic the human brain’s visual representational patterns across object categories and different modalities. Our findings suggest that ReAlnets better align artificial neural networks with human brain representations, making it more similar to human brain processing than traditional computer vision models, which takes an important step toward bridging the gap between artificial and human vision and achieving more brain-like artificial intelligence systems.
@@ 1 Introduction
While current vision models in artificial intelligence (AI) have achieved remarkable advancements, they still fall short of capturing the full complexity and adaptability of the human brain’s information processing. Deep convolutional neural networks (DCNNs) now rival human performance in object recognition [1], and many studies have identified representational similarities between the hierarchical structures of DCNNs and the ventral visual stream [2, 3, 4, 5, 6]. However, aligning DCNNs with human neural representations, though promising, remains an area with significant potential for further exploration. Enhancing the similarity between visual models and the human brain has become a critical concern for both computer scientists and neuroscientists. From a computer vision perspective, brain-inspired models often exhibit greater robustness and generalization, which are essential for achieving brain-like intelligence. Meanwhile, from a cognitive neuroscience perspective, models that more closely resemble brain representations can provide valuable insights into the mechanisms of human visual processing.
@@
Conventional approaches - increasing model depth and layer count - have struggled to emulate the complexity of human visual processing [7]. Researchers have proposed various bio-inspired strategies to leverage the understanding of the human brain to enhance current AI vision models, including altering the architecture of the model (adding recurrent structures [8, 4, 9, 10, 11], dual-pathway models [12, 13, 14, 15, 16], topographic constraints [17, 18, 19, 20] or feedback pathways [21] ) and changing the training task (using self-supervised training [22, 23] or 3D task models [24]). However, an important question remains: Can we directly use human neural activity to align ANNs in object recognition, thereby achieving more human brain-like vision models?
@@
Several previous studies have begun to explore the integration of neural data into machine learning, particularly deep learning models, to enable these models to learn biologically inspired representations from neural data. One common strategy is to introduce a similarity loss during training to increase the representational alignment between models and neural activity, often derived from mouse V1 or monkey V1 and IT regions [25, 26, 27, 28]. Another strategy from [29] integrates an additional task that uses an encoding module to predict monkey V1 neural activity. Both similarity-based methods and multi-task frameworks have been shown to produce more brain-like representations and improve model robustness. However, the key challenge of these neural alignment studies is they can only do single brain region or single model layer alignment. Previous studies typically focused on aligning only a single layer of a CNN with a specific brain region, such as V1 or IT, without a clear understanding of how multiple layers correspond to different brain regions. This oversimplification can lead to misalignment and inaccuracies. Moreover, most of these studies depend on invasive neural recordings from animals instead of human neural data, which limits the direct applicability of findings to human visual processing. Human non-invasive recordings, such as fMRI and EEG, often have lower data quality compared to invasive animal recordings, making it more challenging for models to learn human brain representations effectively. Early attempts involved applying human fMRI signals as additional inputs to machine learning classifiers such as SVMs and CNNs, resulting in improved category classification performance without altering the internal feature representations of the models themselves [30, 31, 32]. While more recent research has directly optimized CNN internal representations align with human fMRI data for video emotion recognition [33], going beyond earlier approaches that only incorporated fMRI features into classification tasks [30], this approach has been limited to a relatively simple six-category emotion classification task. It remains unclear whether it could scale effectively to more complex domains like object recognition, which involves a far greater diversity of categories (e.g., 1000 classes in ImageNet-trained models). Also, it is unclear whether we can apply human neural data to optimize ANNs to achieve more human brain-like internal model representations.
To address these limitations, our study proposes a novel approach that employs an encoding-based framework for multi-layer alignment. This framework goes beyond simple similarity by training an additional encoding module to predict human neural activity, thereby enabling the model to autonomously extract complex visual features. Our approach leverages human neural data to achieve more effective alignment with human brain representations in object recognition tasks.
To bridge the gap between AI vision and human vision, we introduce Re(presentational)Al(ignment)net framework (hereafter, the ReAlnet framework), a method for effectively aligning vision models with human brain representations obtained from non-invasive EEG recordings. EEG was chosen for its high temporal resolution and the capacity to collect a large number of trials through rapid successive stimulus presentation, making it a cost-effective and scalable option. Our novel encoding-based multi-layer alignment framework effectively allows neural networks to learn human brain representations, enabling the creation of personalized vision models tailored to individual neural data. In this ReAlnet framework, each individualized ReAlnet refers to one EEG-aligned model instance trained on a single subject’s data, while the plural ReAlnets refers to the set of ten such individualized models. Here, we define ”direct alignment” as modifying a model’s internal representational structure via a learning objective that explicitly incorporates human neural data. Under this definition, our approach constitutes the first direct alignment of object recognition models using non-invasive human EEG signals. This novel approach opens new possibilities for enhancing brain-like representations in AI models. Furthermore, human EEG-optimized ReAlnets demonstrate improved alignment with human brain representations across different modalities (both human EEG and fMRI) and human behaviors.
consistently show the highest similarity to the human brain. Lines and shading reflect mean±SEM.
@@ 2.1 Aligning CORnet with human EEG representations
In this study, we build and test a novel image-to-brain multi-layer encoding alignment framework (Figure 1). Our core aim is to investigate whether aligning the model with individual neural representations of humans can enhance the model’s similarity to the human brain. As an intial fidelity check, we confirmed that the approach of jointly minimizing classfication and generation losses during the training (Figure 1A, see Methods) yields accurate classification of object category (average Top-1 and Top-5 accuracy on ImageNet: 69.31% and 88.88%, Figure S1A-B) and generates realistic EEG signals (77.85% similarity, computed as Spearman correlation between generated and real EEG time series, treating the full channel × timepoint matrix of each image as a vector, averaged across the test set, Figure S1C-D). While the latter demonstrates that the model successfully learned to reproduce realistic EEG activity, building the most accurate EEG generator is not our main objective. Instead, we seek to use EEG generation as a training mechanism to optimize the model’s internal representations, with the goal of making them more similar to human brain representations. Our critical hypthesis is that modifying the model’s internal representations in this way may carry benfits that generalize across image diet and modalities. In the following sections, we evaluate whether the model’s internal representations have indeed become more brain-like, by conducting representational similarity analysis (RSA to test models’ similarity to human EEG, human fMRI, and behavior.
@@
To apply this alignment framework, we built ten individual ReAlnets, using the state-of-the-art CORnet-S model [34, 9] as the foundational architecture. Each ReAlnet, which has the same architecture as CORnet, is additionally trained on a real human subject’s EEG signals, recorded while viewing a massive number of natural images from THINGS EEG2 [35] training set. After training, we employed an independent test dataset consisting of 200 images and associated EEG activity from the THINGS EEG2 test set. These test set images had not been presented at all during the training process, coming from entirely novel (untrained) object categories. We input these 200 test images to each model and obtained the feature vectors corresponding to each image for each layer in the model. Then we calculated the temporal similarity between different models and human brain EEG using RSA method. Figure 1B shows the representational similarity between different models and temporal human brain EEG signals for early and late layers. Comparing the ReAlnets to the original CORnet-S, along with ResNet101 and CLIP (with a ResNet101 backbone) as additional baselines, ReAlnets consistently show the highest similarity to the human brain patterns.
@@
Our assessment of the models’ similarity to humans is not limited to its similarity with human brain EEG representations (based on THINGS EEG2 test set [35]); Similar to EEG, we then evaluated the model’s similarity to human brain fMRI representations (a completely different modality) from human subjects viewing novel image categories (based on Shen fMRI test set [36]). Additionally, we measured the similarity between the model and human behavior in several object recognition tasks using the Brain-Score platform [37] based on two behavioral benchmarks. For methodological details, refer to our Methods section. In the following sections, we elaborate on how ReAlnets – compared with the original CORnet not aligned with human neural data – more closely resembles human brain representations.
@@ 2.2 Improved similarity to human EEG
Here, for each of the 10 human subjects from the THINGS EEG2 dataset [35], we calculated (1) the similarity between their EEG data and CORnet (with the same structure as ReAlnets, but non aligned human neural data and non-individualized model), and (2) the similarity between their EEG data and the subject-matched ReAlnet, via RSA based on THINGS EEG2 test dataset (See more details in Methods section). As an additional control, we trained a Scrambled-model version of the ReAlnet framework by aligning on 10 subjects’ scrambled EEG signals (more details in Control Experiments section and Methods section) and conducted the same RSA analysis.
ReAlnets exhibit significantly higher similarity to human EEG neural dynamics for all four visual layers (Layer V1: 70-130ms and 160-200ms; Layer V2: 60-200ms; Layer V4: 60-200ms; Layer IT: 70-160ms) than the original CORnet without human neural alignment, or than Scrambled models trained on scrambled EEG signals (Figure 2A). The EEG similarity curves often show a two-peaked shape, potentially reflecting different stages of visual information processing. Based on prior studies on human neural dynamic representations [38, 39, 40], the earlier peak (1̃00ms) likely corresponds to the processing of lower-level visual features, such as color and retinal size. And the later peak (1̃50-200ms or even later) may reflect the processing of higher-level semantic features, such as real-world size and animacy information.
Further statistical analysis of each layer’s similarity improvement (ReAlnet - CORnet) and improvement ratio ((ReAlnet - CORnet) / CORnet) also indicates that at the similarity peak timepoint, there is a maximum of a 6% similarity improvement and a 40% improvement ratio (Figure 2B). Importantly, all indiviudalized ReAlnets showed positive improvement over CORnet, suggesting that our ReAlnet framework robustly generalizes across subjects that effectively enhances model-EEG similarity using each individual’s EEG data. It is worth noting that the test set doesn’t overlap with the training set in terms of object categories (concepts). Therefore, these significant improvements reveal ReAlnets’ generalization capability across different object categories. In addition to report layer-wise RSA curves, we also computed the maximum similarity across all layers for each model at each time point to provide a clearer summary of model-EEG alignment (Figure 2C).
@@
Additionally, unlike traditional models in computer vision, ReAlnet is a personalized model trained based on different individuals’ neural data. Just as neuroscientists are interested in studying individual differences in the brain, we are similarly interested in exploring individual differences across the ten personalized ReAlnets - whether ReAlnets exhibit intra-model individual variabilities and how such variabilities change across different layers of the model. Note that CORnet serves as a single, publicly available reference model with fixed official weights, and thus does not have individualized versions analogous to ReAlnets. Therefore, inter-model variability can only be meaningfully assessed across individualized ReAlnets. Importantly, all ten ReAlnets were initialized with the same pretrained CORnet-S weights and trained using the same random seed, ensuring that any observed differences arise solely from the individual EEG data used for fine-tuning, not from variations in model initialization. We hypothesize that the observed individual differences in ReAlnets may provide insights into potential mechanisms underlying individual differences in human brains. To investigate this, we also analyzed individual variability in human brain regions (V1, V2, V4, and LOC) using the Shen fMRI dataset based on a similar RDM-based correlation analysis. As a comparison, we also conducted comparisons between model RDMs based on 200 images in THINGS EEG2 test set across different layers, using the dissimilarity (one minus the Spearman correlation coefficient) between two RDMs corresponding to two ReAlnets as an individual variability index. Importantly, we included the output layer of ReAlnets, which performs category classification, and found that its individual variability decreases, mirroring the pattern seen in the LOC region of the human brain (additional analysis in Figure S2 confirms that LOC is significantly more similar to the model’s output layer than Layer IT, suggesting that the model’s output layer rather than Layer IT is more comparable to brain LOC region). While humans also show hierarchical increases in between-subject variability, the larger absolute scale may partly arise from cross-individual measurement noise in neuroimaging. Our results suggest: (1) Personalized ReAlnets indeed exhibit individual variability (Figure 2D). (2) This variability increases with the depth of the layers (from Layer V1 to Layer IT, Figure 2D) and decreases in the output layer, which aligns with the trend observed in the human brain - where individual variability increases from V1 to V4 and then decreases in LOC.
To further evaluate the generalization ability of individualized ReAlnets within the EEG modality, we conducted a cross-subject analysis, testing each model against the EEG RDMs of all subjects. We computed a model–subject similarity matrix where each entry represents the average RSA similarity across four model layers and the 50-200 ms time window. This approach summarizes each ReAlnet’s representational match with every subject and produces a compact confusion matrix (Figure 2E). To more directly assess cross-subject generalization, we subtracted each subject’s similarity with CORnet from all cells in that subject’s column (Figure 2F). Positive values in this normalized matrix indicate that even mismatched ReAlnets outperform the CORnet baseline, confirming that ReAlnets capture representational structures that generalize across individuals. Finally, to control for differences in signal quality across subjects, we applied column-wise normalization to the matrix (Figure 2G). To more directly quantify subject-specificity, we compared similarity values for matched model-subject pairs (the diagonal: i.e., models evaluated on the same subject used for training) against unmatched pairs (the off-diagonal) and observed that matched model–subject pairs show significantly higher similarity than unmatched pairs (=5.6068, =.0003). These results confirm that individualized ReAlnets capture subject-specific representational features, and that these models generalize across individuals within the same neural modality.
@@ 2.3 Improved similarity in ReAlnets to human fMRI
Although ReAlnets demonstrate higher similarity to human EEG, a question arises: Do ReAlnets learn representations specific to EEG, or more general neural representations of the human brain? To ensure that our alignment framework enables the model to learn representations beyond the single modality of EEG, we utilized additional human fMRI data of three human subjects viewing natural images to evaluate the model’s cross-modality representational similarity to human fMRI.
Excitingly, we indeed observed a clear increase in this cross-modal brain-like similarity. Based on human fMRI signals of three subjects viewing 50 natural images, the similarity results indicate that, overall, ReAlnets exhibit a stronger resemblance to human fMRI data than CORnet and Scrambled models, despite being aligned using a different type of neural signal with very different spatiotemporal properties (human EEG data), and that was collected from a different set of participants (Figure 3A). Although a few conditions show CORnet may have higher similarity than ReAlnets, the general trend favors ReAlnets - highlighting their stronger alignment with human neural representations and showing that EEG-optimized ReAlnets generalize broadly to human representations. Moreover, the similarity improvements across modalities were significantly correlated (Figure S3): the ReAlnets showing more similarity improvement when tested on EEG data also showed more similarity improvement when tested on fMRI data, (), indicating that the ReAlnet alignment framework is tapping into meaningful and robust human neural signals.
@@
Additionally, to further assess the enhanced similarity between ReAlnets and human brain representations during image observation - not solely with natural images - we also measure the similarity between models and human fMRI signals from three subjects viewing 40 artificial shape and 10 alphabetical letter images in the Shen fMRI test dataset. Although these images might be outside the natural image distribution, our results further demonstrate ReAlnets’ improved similarity to human brain representations in comparisons to CORnet and Scrambled models (Figure 3B-C). While some comparisons yield small absolute differences, we also examined effect sizes to better assess their statistical and practical significance (detailed statistical results including -value, -value, and Cohen’s are listed in Table S1). The stronger effects ( and —Cohen’s —¿1.0; bolded in Table S1) were observed in brain region V2 and V4 under natural images condition and in brain region V1 and V2 under artifical shape images condition, indicating particularly robust representational improvements in these regions under different conditions. Other ROIs and conditions also showed statistically significant improvements, though with relatively smaller effect sizes, suggesting that the enhancement is widespread but especially pronounced in certain areas.
@@
These findings further highlight three points: (1) Across multiple ROIs, ReAlnets exhibits higher human fMRI similarity than CORnet. (2) Despite being trained with the EEG data of subjects not in the fMRI dataset, almost every ReAlnet shows higher fMRI similarity, suggesting that ReAlnets learn consistent brain information processing patterns across subjects. (3) Images from the fMRI datasets for evaluation were never presented during the alignment training, reaffirming the generalization of ReAlnets in improving brain-like similarity across object categories and images.
2.4 Improved similarity in ReAlnets to behavior